OpenStudy (anonymous):

How do you simplify: (2x^2+x-12) /(x^2-16)

OpenStudy (anonymous):

I was trying to factor the top to see if anything in the top and bottom cancel but i'm not sure.

hero (hero):

Okay

hero (hero):

You have to use long division for this

OpenStudy (anonymous):

im not sure how you do long divison with expressions

hero (hero):

Hold on a second

hero (hero):

Actually, that is the wrong solution. It shouldn't be set to zero

hero (hero):

Are you sure you have this written correctly? Double check

OpenStudy (anonymous):

Hero)))

OpenStudy (anonymous):

sure yeah its correct: (2x^2 + x -12 )/ (x^2 - 16)

OpenStudy (anonymous):

I send u an email)

OpenStudy (anonymous):

to the gmail)))

hero (hero):

Well, what do you have to do? Solve for x?

OpenStudy (anonymous):

the question was just to simplify the expression

OpenStudy (anonymous):

so i was trying to see if anything cancels in teh top or bottom

hero (hero):

\[\frac{2x^2 +x - 12}{(x+4)(x-4)}\]

hero (hero):

That's the best you can do

hero (hero):

Okay Kina

OpenStudy (anonymous):

i got up to that part...so there is nothing aafter taht?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

ok tahnks...can you help me with another simplify teh expression question?

OpenStudy (anonymous):

If I can, lol.. sure.

OpenStudy (anonymous):

thanx...Simplify : (1/x+h) - (1/x)

OpenStudy (anonymous):

Make them have the same denominator.

OpenStudy (anonymous):

so is teh lcd: x+h

OpenStudy (anonymous):

No. You need to multiply the denominators, right? It would be x(x+h). Remember whatever you do to the bottom, you have to do to the top...

OpenStudy (anonymous):

oh ok...lemme try it out

OpenStudy (anonymous):

After I multiply teh top and bottom out i get (x-x+h) / (x+h)...so does that simplify into h/ x+h?

OpenStudy (anonymous):

Hero, I got -h/x(x+h) ^^

hero (hero):

yeah, that's right

hero (hero):

Good job

hero (hero):

I forgot the negative

OpenStudy (anonymous):

@denebelbel can u show me how u got ur solution?

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ x+h\end{matrix}\right) - \left(\begin{matrix}1 \\ x\end{matrix}\right) = \left(\begin{matrix}x-(x+h) \\ x(x+h)\end{matrix}\right) = \left(\begin{matrix}x-x-h \\ x(x+h)\end{matrix}\right) = \left(\begin{matrix}-h \\ x(x+h)\end{matrix}\right)\]

hero (hero):

I helped you with one of your other problems. No medal? Not fair, lol

OpenStudy (anonymous):

I ll give both of you a medal =]

hero (hero):

Okay

OpenStudy (anonymous):

@ deneble and @hero tahnk u both for helping...i suck at math XD